OpenStudy (anonymous):

write the expresson as a single logarithm. 1/2(x^2+1)-4ln(1/2)-(1/2)[ln(x-4)+ln(x)]

OpenStudy (anonymous):

do you have ln in front of the first term ?

OpenStudy (anonymous):

ya sorry it it 1/2ln then the rest

OpenStudy (anonymous):

k

OpenStudy (anonymous):

1st term ln(x^2+1)^1/2

OpenStudy (anonymous):

2nd term ln(1/2)^4

OpenStudy (anonymous):

3rd term ln(x^2-4x)^1/2

OpenStudy (anonymous):

got that far anf then lost

OpenStudy (anonymous):

ln a +ln b = ln(ab) and ln a - ln b = ln(a/b)

OpenStudy (anonymous):

can you do this now

OpenStudy (anonymous):

so my problem is all - so the 1st and 2nd term divided then all that divded by the 3rd term

OpenStudy (anonymous):

yes...left to right

OpenStudy (anonymous):